• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 12:14:00 on October 22, 2009

    Entry number 296100

    keyword=Spectrometer magnet problems diagnosed, being fixed

    The valve settings for Q1 and Q2 were off by 25%, which explains the instability that was preventing them from being turned on. Ed and his crew are working on recalibrating the valves. The magnet problems should be fixed by today 15:00. This will put us one hour behind MCC's expected turn-on time, but hopefully with stable magnet conditions.